Bond Valuation with Annual Payments Jackson Corporation's bonds have 5 years remaining to maturity. Interest is paid annually, the bonds have a $1,000 par value, and the coupon interest rate is 8.5%. The bonds have a yield to maturity of 11%. What is the current market price of these bonds? Round your answer to the nearest cent.?
Answer #1

Annual coupon=1000*8.5%=85

Hence current price=Annual coupon*Present value of annuity factor(11%,5)+$1000*Present value of discounting factor(11%,5)

=85*3.69589702+1000*0.593451328

=$907.60(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=85[1-(1.11)^-5]/0.11

=85*3.69589702

2.Present value of discounting factor=1000/1.11^5

=1000*0.593451328